# Back from the dead: AT2019aalc as a candidate repeating TDE in an AGN

### v2

@@ 5.1.3 Coronal lines
Based on a study of SDSS spectra, Wang et al. (2012) reported the discovery of Extreme Coronal Line Emitter (ECLE) galaxies which show extremely strong coronal lines. Three TDEs (AT2017gge, AT2019qiz and AT2022upj) were also found to exhibit coronal lines (Onori et al. 2022; Newsome et al. 2022; Short et al. 2023). The BFF AT2021loi (Makrygianni et al. 2023) and the spectacular transient event AT2019avd (Malyali et al. 2021) also exhibit strong coronal lines. The BFF F01004-2237 showed variable coronal lines on a timescale of years. Based on the publicly available spectra of AT2017bgt on the Transient Name Server777https://www.wis-tns.org/ (TNS) we see enhanced coronal line emission in this BFF as well.
@@
Wang et al. (2012) state that the most possible origin of the ECLEs are remnants of an earlier TDE outburst which was later confirmed by Short et al. (2023) who revealed that the TDE AT2019qiz more likely follows the line ratio correlations of ECLEs instead of regular AGN. The coronal lines in AT2019qiz appeared around days after the optical peak, further supporting this scenario. Moreover, the estimated rate of coronal line emitters is consistent with the estimated rate of TDEs (Wang et al. 2012). Later, Callow et al. (2024) discovered 9 new ECLEs using SDSS spectra and refined the previous rate calculations. They found that the ECLE rates are below the TDE rates, however, can still be consistent with assuming that only to per cent of all TDEs produce variable coronal lines. According to Wang et al. (2012), the origin of the coronal lines in ECLEs can be explained with the liberation of Fe from dust grains that were destroyed by the TDE flare. A follow-up spectroscopic analysis of the original ECLE sample of Wang et al. (2012) revealed that the coronal line signatures did not recur in five out of the seven objects, supporting their transient light echo origin (Clark et al. 2024). The large dust echo of AT2019aalc is consistent with this picture. Furthermore, the TDEs with coronal line detection have more luminous IR dust echo compared to other optical TDEs. AT2019qiz is one of the few TDEs included in the accretion flare sample of van Velzen et al. (2024) due to its strong dust echo. Another coronal line-detected TDE AT2017gge is part of the mid-infrared (MIR) dust echo flare sample of Hinkle (2024) which contains ambiguous nuclear transients with high dust covering factors. The extreme dust echo seems to characterize the BFFs as well. The BFFs AT2021loi, AT2017bgt, OGLE17aaj and the candidate BFF AT2019avd are present in the above-mentioned MIR dust echo sample (Hinkle 2024) and of them are also part of the Flaires sample of dust-echo-like IR flares (Necker et al. 2024). The Bowen line transient AT2019pev shows strong high-ionization coronal lines (Frederick et al. 2021; Yu et al. 2022) and is also part of the MIR flare sample of Hinkle (2024). These findings suggest a connection not only between the ECLEs and TDEs, but also the BFFs. We performed modeling of the publicly available LRIS spectrum of AT2017bgt taken on 2017-02-23 using PyQSOFit, following the procedure described in Subsect. 3.5. Fig. 13 shows the [Fe x] versus the [O iii luminosities of the ECLEs studied by Wang et al. (2012), AT2019qiz (Short et al. 2023), two BFFs and AT2019aalc together with the SDSS Seyfert sample of Gelbord et al. (2009). This clearly implies an offset of the ECLEs, BFFs and AT2019aalc from the regular AGN sample.
@@ 5.2 Optical re-brightening
The long-term ZTF light curve of AT2019aalc shows two well-separated flares. Both flares evolved quickly and reached their peaks within months, followed by a slow decay on yearly timescales. Generally, TDEs are characterized by smoothly and rapidly declining light curves () as expected from mass fallback considerations (Rees 1988) without any prominent bumps in the case of full disruption. AT2019aalc in turn has a slowly decaying first flare with a power-law fit returning a power-law index of and a second flare that began years after the initial flare. The first systematically identified repeating partial TDE (AT2020vdq), however, shows not just a double-peaked long-term optical light curve, but its second flare is times more luminous compared to the first one (Somalwar et al. 2023). This implies an even more significant optical re-brightening than the one detected in AT2019aalc. The second flare of partial TDEs is expected to follow a steep declining with a power-law of . In the case of AT2020vdq, the second flare declined much slower (Somalwar et al. 2023) than this but still clearly faster than the second flare of AT2019aalc which declined with a power-law index of .
@@
Interestingly, most of the sources of the small sample of classified BFFs show optical re-brightening or bumps. Fig. 14 shows the optical evolution of the three BFFs with detected re-brightening episodes together with AT2019aalc. The optical light curve of AT2017bgt shows a bump roughly days after its initial flare which is also clearly visible in the binned light curve published in Trakhtenbrot et al. (2019). Originally, F01004-2237 had undergone a luminous flare (Tadhunter et al. 2021) in 2010 and later Trakhtenbrot et al. (2019) classified it as a BFF. The binned light curve of F01004-2237 published in Makrygianni et al. (2023) shows a bump around a year after its peak. We investigated the recent activity of this source and found that it significantly re-brightened in based on its ATLAS monitoring. The recent flare was more luminous than the first one and declined faster. In addition to these two BFFs, AT2021loi shows a bump when declining, roughly a year after the first flare reached its peak. Makrygianni et al. (2023) suggests that optical re-brightening or bumps seen when decaying might be common features of BFFs. Only one transient classified as a BFF, OGLE17aaj, has not shown signs of optical re-brightening or bumps in its light curve to date. The first flare of the studied BFFs decay on yearly timescales, much slower than regular TDE flares but more similar to AT2019aalc. However, the BFFs with re-brightening episodes tend to have a faster decaying second flare/bump in comparison with their initial flare, while in the case of AT2019aalc the second flare decays slower. The BFFs classified so far tend to have remarkable variances in their optical light curve evolution despite the similarities we see. These might be explained by different subclasses of BFFs, similarly to regular TDEs, where different subclasses were defined based on their light curve evolution (e.g., Charalampopoulos et al. 2023).
@@ 5.3.1 Possible explanations of the long-term radio flare
In this subsection we discuss the long-term radio variability of AT2019aalc. The source experienced a brightening of its flux density by a factor of on a timescale of less than years followed by a long-term plateau ( years) indicated by the archival radio survey data and our ATCA monitoring. Although radio variability on yearly timescales is a common phenomenon of Seyfert galaxies, the sharp, high-amplitude radio flare and especially the long-term plateau are generally not seen in intrinsically variable Seyfert galaxies (e.g., Mundell et al. 2009; Koay et al. 2016) and therefore require further investigation. The unusual radio spectrum with a high frequency turnover is also discussed below.
@@
The brightness temperature of K at GHz significantly exceeds K, known as an upper limit for supernova remnants and He ii regions in star-forming galaxies (Condon 1992). Consequently, the radio emission we detected must be non-thermal and originate from physical processes associated to AGN-like (including TDEs) activity (see Bontempi et al. 2012 as an example how the estimated brightness temperatures help to characterize the radio emission of Seyfert galaxies). This way, we can also safely rule out thermal free-free emission from the accretion disk or the dusty torus (e.g., the case of NGC 1068, Gallimore et al. 1997) as the origin of the detected radio emission.
@@
Advection-dominated accretion flows have also been proposed to explain the origin of the radio emission of low-luminosity AGN (e.g., Seyfert galaxies, Mahadevan 1997). Assuming that accretion flows dominate the radio emission of AT2019aalc, following Yi & Boughn (1998) we estimate an expected GHz radio luminosity of erg s-1, a magnitude of lower than the observed one, clearly indicating that these cannot explain the observed radio emission.
@@
Magnetized coronal winds originating from the AGN accretion disk could produce the non-thermal radio emission of AT2019aalc. Laor & Behar (2008) found a correlation of studying the Palomar–Green (PG: Green et al. 1986) radio-quiet quasar sample (in good agreement with the correlation found for coronally active stars by Guedel & Benz (1993). Later, Behar et al. (2015) came to the same conclusion using a sample of seven radio-quiet Seyfert galaxies. For our source, based on the mean ATCA C-band and the extrapolated Swift/XRT keV luminosities, we estimate a ratio of suggesting that magnetized coronal winds alone cannot fully explain the observed radio emission of AT2019aalc. Moreover, coronal emission should be highly variable at radio wavelengths.
@@
Out-flowing synchrotron emitting material induced by enhanced accretion are commonly detected in AGN and TDEs in the forms of non-relativistic outflows or relativistic jets (e.g., Paragi et al. 2017; Mohan et al. 2022; Kozák et al. 2024). The high brightness temperature of the EVN-detected radio feature indicates its radio core identification. The steep () radio spectrum indicate the presence of another, steep spectrum synchrotron-emitting component(s) in the system as radio cores are characterized by a flat radio spectrum. The spectrum may be the composition of the flat spectrum radio core and the additional, steep spectrum component(s). The high-frequency turnover around GHz supports the idea of a newly ejected radio-emitting component.
@@ 5.3.2 The radio spectrum
The turnover suggests the multiple component nature of the spectrum. Radio spectra with high-frequency excess have been observed in a few radio-quiet narrow-line Seyfert 1 galaxies (Berton et al. 2020). These sources are thought to be kinematically young AGN with small linear sizes or heavily absorbed AGN. The high-frequency peak might be explained with the kinematic age of a newly ejected component. These extremely young radio sources therefore peak at very high frequencies (e.g., Berton et al. 2020). The low-frequency peak is, however, more likely to be explained with star formation activity instead of AGN emission since the latter is absorbed due to synchrotron self-absorption (SSA) or free-free absorption (FFA).
However, in the case of AT2019aalc, the large-amplitude variability concluded from the VLASS observations and the EVN detection of a high brightness temperature ( K) feature at GHz clearly imply AGN-like activity over a star formation scenario at these frequencies and is incompatible with significant absorption caused by SSA or FFA. Therefore, our results indicate that the radio spectrum of AT2019aalc is strongly dominated by AGN activity below GHz as well, even if a contribution from star formation, especially at GHz, is probable.
A newly ejected component can explain the high-frequency excess whilst the low-frequency part is originating from optically thin synchrotron emission of the AGN. Thus, the overall spectrum represents a superposition of past and recent AGN activity, indicative of its recently enhanced, intermittent behavior. Järvelä et al. (2021) states that restarted AGN activity can explain at least some of the radio spectra studied by Berton et al. (2020). Intermittent activity is further supported by theoretical models highlighting that sources with high accretion rates are more prone to this kind of behavior (Czerny et al. 2009).
@@ 5.4 X-ray properties
The X-ray spectra of AGN at energies above keV have a power-law-like shape. These X-ray photons are believed to be triggered by Compton up-scattering of the accretion disk photons off hot electrons surrounding the disk, in a hot ( K) optically thin corona above the disk (e.g., Kammoun et al. 2015). Below keV energies many AGNs show an excess in their spectrum referred to as the soft X-ray excess. It can be well modeled by a blackbody model with a best-fit temperature in the range keV. The origin of the excess might be explained via Comptonized disk emission and the reflection of hard X-ray photons from the surface of the disk can account also for the excess (Done et al. 2012).
The X-ray spectrum of AT2019aalc exhibits a non-thermal component with a power-law index of and an unusually soft thermal component with a blackbody temperature of eV. Most of the BLSy1s typically have a harder X-ray spectrum with a power-law index of (e.g., Pons & Watson 2014). Our value is more commonly seen in sources characterized by enhanced accretion, e.g., flaring NLSy1s or TDEs which have typically soft X-ray spectra with (Boller et al. 1996; Saxton et al. 2020).
The soft X-ray flux of Seyfert galaxies was found to be well correlated with the observed fluxes in the coronal line [Fe x]. The coronal lines to X-ray flux ratio is log for both broad and narrow-line Seyfert 1 galaxies (Gelbord et al. 2009). For AT2019aalc we estimate values of log when optical spectroscopic and X-ray observations were performed nearly simultaneously (due to the rapid X-ray variability, we consider only observations taken on the same day). Thus, the soft X-ray photons excite the coronal lines continuously and not years later originating from a past TDE as suggested for ECLEs by Wang et al. (2012).
The reoccurring optical/UV bumps seen during the decaying of the main optical flares are likely driven by the soft X-ray flares which lead these outbursts only with a couple of weeks. The soft X-ray excess appears to be more dominant during these flares. Interestingly, the soft X-ray flares show hints of quasi periodicity of days. A few Repeating Nuclear Transients (RNTs) were discovered in the past few years: Quasi-Periodic Eruptions (QPEs) exhibit with soft X-ray flares on time-scales of hours to days while partial TDEs on timescales of months to years (e.g., Chakraborty et al. 2021; Quintin et al. 2023; Miniutti et al. 2023; Nicholl et al. 2024). We note that another BFF, AT2020afhd, shows hints of quasi-periodic soft X-ray flares according to its Swift/XRT monitoring. This source is included in the The Living Swift XRT Point Source Catalogue (LSXPS: Evans et al. 2023) as LSXPS J031335.6020905. The potential connection between BFFs and QPEs should be explored in future studies.

### v3

@@ 5.1.3 Coronal lines
Based on a study of SDSS spectra, 2012ApJ...749..115W reported the discovery of seven Extreme Coronal Line Emitter (ECLE) galaxies, which show extremely strong coronal lines. Numerous TDEs were also found to exhibit coronal lines. The BFF AT2021loi (AT2021loi) and the spectacular transient event AT2019avd (avd1) also exhibit strong coronal lines. The BFF F01004-2237 showed variable coronal lines on a timescale of years (f01). Based on the publicly available spectra of AT2017bgt on the Transient Name Server777https://www.wis-tns.org/ (TNS) we detect enhanced coronal line emission in this BFF, too.
@@
The most plausible origin of the ECLEs is that they represent echoes of a past TDE outburst (2012ApJ...749..115W; 2024MNRAS.528.4775H; 2024MNRAS.532..112K; 2024ApJ...977..258N). Notably, 2023MNRAS.525.1568S revealed that the TDE AT2019qiz more likely follows the line ratio correlations of ECLEs instead of regular AGN. The coronal lines in AT2019qiz appeared around days after the optical peak, further supporting this scenario. Moreover, the estimated rate of coronal line emitters is consistent with the estimated rate of TDEs (2012ApJ...749..115W). Later, ecle_rate_callow discovered 9 new ECLEs using SDSS spectra and refined the previous rate calculations. They found that the ECLE rates are below the TDE rates, however, can still be consistent with the assumption that only to percent of all TDEs produce variable coronal lines. According to 2012ApJ...749..115W, the origin of the coronal lines in ECLEs can be explained with the liberation of Fe from dust grains that were destroyed by the echo of past TDE flares. A follow-up spectroscopic analysis of the original ECLE sample of 2012ApJ...749..115W revealed that the coronal line signatures did not recur in five out of the seven objects, supporting their transient light echo origin (ecle_mw). The large dust echo of AT2019aalc is consistent with this picture. Furthermore, the TDEs with coronal line detection have more luminous IR dust echoes compared to other optical TDEs. AT2019qiz is one of the few TDEs included in the accretion flare sample of Lancel, due to its strong dust echo. Another coronal line-detected TDE AT2017gge is part of the mid-infrared (MIR) dust echo flare sample of 2024MNRAS.531.2603H, which contains ANTs with high dust covering factors. The extreme dust echo seems to characterize the BFFs as well. The BFFs AT2021loi, AT2017bgt, OGLE17aaj and the candidate BFF AT2019avd are present in the above-mentioned MIR dust echo sample (2024MNRAS.531.2603H) and two of them are also part of the Flaires sample of dust-echo-like IR flares (2025A&A...695A.228N). The BFF F01004-2237 exhibited a MIR flare (2017ApJ...841L...8D). The Bowen line transient AT2019pev shows strong high-ionization coronal lines (2021ApJ...920...56F; 2022MNRAS.515.5198Y) and is also part of the MIR flare sample of 2024MNRAS.531.2603H. These findings suggest a connection not only between the ECLEs and TDEs, but also the BFFs.
Fig. 13 shows the [Fe x] versus the [O iii luminosities of the ECLEs, the coronal-line detected TDEs and BFFs and AT2019aalc, together with the coronal-line detected SDSS Seyfert sample of 2009MNRAS.397..172G. For this plot, we performed modeling of the publicly available spectra888We retrieved the LRIS spectra from the Transient Name Server (TNS; https://www.wis-tns.org) and the ESO Faint Object Spectrograph and Camera version 2 (EFOSC2) spectra from the ESO Science Portal (archive.eso.org/scienceportal) of several TDEs and BFFs using PyQSOFit, following the procedure described in Subsect. 3.5. These are the TDEs: AT2017gge (2017-09-14, EFOSC2); AT2018dyk (2018-08-08, LRIS); AT2018bcb (2018-05-06, EFOSC2); AT2021qth (2022-05-26, LRIS); AT2022upj (2022-10-26, DBSP); and the BFFs: AT2017bgt (2017-02-23, LRIS) and F01004-2237 (2018-08-13, EFOSC2). Absolute fluxes, and hence derived luminosities, are subject to systematic uncertainties from slit losses and atmospheric extinction differences. We adopt a conservative % uncertainty on the flux scale to account for these effects, which is combined in quadrature with the statistical uncertainties from the PyQSOFit spectral fits. This method is applied only to cases where the slit position angle differed at least five degrees from the parallactic angle. Fig. 13 clearly implies an offset of the ECLEs, BFFs and AT2019aalc from the regular AGN sample. We note that the only object in the sample of 2009MNRAS.397..172G classified as a galaxy rather than an AGN is, in fact, the same source as one of the variable ECLEs (which appear as nearly overlapping points in our plot). No publicly available spectra were found of the TDEs AT2021dms, AT2024mvz and AT2021acak while in the case of AT2017gge, it was not possible to safely distinguish [Fe x] from H. We note that AT2021acak exhibited two distinct optical flares (2023RAA....23b5012L), with the second being more luminous, and showed a He ii N iii complex similar to that of AT2019aalc.
@@ 5.2 Optical re-brightening
The long-term ZTF light curve of AT2019aalc shows two well-separated flares. Both flares rose quickly and reached their peaks within two months, followed by a slow decay on yearly timescales. Generally, TDEs are characterized by smoothly and rapidly declining light curves () as expected from mass fallback considerations (1988Natur.333..523R), and without any prominent bumps in the case of full disruption. AT2019aalc in turn has a slowly decaying first flare with a power-law fit returning a power-law index of and a second flare that began four years after the initial flare. The first systematically identified repeating partial TDE (AT2020vdq), however, shows not just a double-peaked long-term optical light curve, but its second flare is times more luminous compared to the first one (2020vdq). This implies an even more significant optical re-brightening than the one detected in AT2019aalc. In contrast, the candidate pTDE AT2022dbl exhibits with a brighter first flare. The second flare of partial TDEs is expected to follow a steep declining with a power-law of . In the case of AT2020vdq and AT2022dbl, the second flares declined slower than this (2020vdq; 2025ApJ...987L..20M), but still faster than the second flare of AT2019aalc, which declined with a power-law index of . Notably, AT2020vdq and AT2022dbl both exhibit slower-declining second flares compared to their first flares, similar to AT2019aalc.
@@
Interestingly, most of the sources of the small sample of classified BFFs show optical re-brightening or bumps. Fig. 14 shows the optical evolution of the three BFFs with detected re-brightening episodes, together with AT2019aalc. The optical light curve of AT2017bgt shows a bump roughly days after its initial flare which is also clearly visible in the binned light curve published in Trakhtenbrot. Originally, F01004-2237 had undergone a luminous flare (f01) in 2010 and later Trakhtenbrot classified it as a BFF. The binned light curve of F01004-2237 published in AT2021loi shows a bump around a year after its peak. We investigated the recent activity of this source and found that it significantly re-brightened in based on its ATLAS monitoring. The recent flare was more luminous than the first one and declined faster. In addition to these two BFFs, AT2021loi shows a bump when declining, roughly a year after the first flare reached its peak. AT2021loi suggests that optical re-brightening or bumps seen when decaying might be common features of BFFs. OGLE17aaj, has not shown a clear optical re-brightening, however a bump likely appears roughly a year after the peak, similar to AT2017bgt and AT2021loi. The BFFs classified so far tend to have remarkable variances in their optical light curve evolution despite the similarities. These might be explained by different subclasses of BFFs, similarly to regular TDEs, where different subclasses were defined based on their light curve evolution (e.g., 2023A&A...673A..95C).
@@ 5.3.1 Possible explanations of the long-term radio flare
In this subsection we discuss the long-term radio variability of AT2019aalc. The radio source experienced a brightening of its flux density by a factor of two on a timescale of less than years, followed by a long-term plateau ( years) indicated by the archival radio survey data and our ATCA monitoring. Although radio variability on yearly timescales is a common phenomenon of Seyfert galaxies, the sharp, high-amplitude radio flare and especially the long-term plateau are generally not seen in intrinsically variable Seyfert galaxies (e.g., seyf_radio_var; 2016MNRAS.460..304K) and therefore require further investigation. The unusual radio spectrum with a high frequency turnover is also discussed below.
@@
The brightness temperature of K at GHz significantly exceeds K, known as an upper limit for supernova remnants and He ii regions in star-forming galaxies (sfrTb). Consequently, the radio emission we detected must be non-thermal and originate from physical processes associated to AGN-like (including TDEs) activity (see 2012MNRAS.426..588B as an example of how the estimated brightness temperatures help to characterize the radio emission of Seyfert galaxies). This way, we can also safely rule out thermal free-free emission from the accretion disk or the dusty torus (e.g., the case of NGC 1068, 1997Natur.388..852G) as the origin of the detected radio emission.
@@
Advection-dominated accretion flows have also been proposed to explain the origin of the radio emission of low-luminosity AGN (e.g., Seyfert galaxies, 1997ApJ...477..585M). Assuming that accretion flows dominate the radio emission of AT2019aalc, following 1998ApJ...499..198Y we estimate an expected GHz radio luminosity of erg s-1, orders of magnitude lower than the observed one, clearly indicating that these cannot explain the observed radio emission.
@@
Magnetized coronal winds originating from the AGN accretion disk could produce the non-thermal radio emission of AT2019aalc. 2008MNRAS.390..847L found a correlation of studying the Palomar–Green (PG: 1986ApJS...61..305G) radio-quiet quasar sample (in good agreement with the correlation found for coronally active stars by 1993ApJ...405L..63G). Later, 2015MNRAS.451..517B came to the same conclusion using a sample of seven radio-quiet Seyfert galaxies. For our source, based on the mean ATCA C-band and the extrapolated Swift/XRT keV luminosities, we estimate a ratio of , suggesting that magnetized coronal winds alone cannot fully explain the observed radio emission of AT2019aalc. Moreover, coronal emission should be highly variable at radio wavelengths.
@@
Out-flowing synchrotron-emitting material induced by enhanced accretion is commonly detected in AGN and TDEs in the forms of non-relativistic outflows or relativistic jets (e.g., 2017IAUS..324..119P; 2022ApJ...927...74M; 2024Galax..12....8K). The high brightness temperature of the EVN-detected radio feature indicates its radio core identification. The steep () radio spectrum indicates the presence of another, steep-spectrum synchrotron-emitting component(s) in the system, as radio cores are characterized by a flat radio spectrum. The spectrum may be the composition of the flat-spectrum radio core and the additional, steep-spectrum component(s). The high-frequency turnover around GHz supports the idea of a newly ejected radio-emitting component.
@@ 5.3.2 The radio spectrum
The turnover suggests the multi-component nature of the spectrum. Radio spectra with high-frequency excesses have been observed in a few radio-quiet narrow-line Seyfert 1 galaxies (2020A&A...636A..64B). These sources are thought to be kinematically young AGN with small linear sizes or heavily absorbed AGN. The high-frequency peak might be explained with the kinematic age of a newly ejected component. These extremely young radio sources therefore peak at very high frequencies (e.g., 2020A&A...636A..64B). The low-frequency peak is more likely originating from star formation activity instead of AGN emission, since the latter is absorbed at these frequencies due to synchrotron self-absorption (SSA) or free-free absorption (FFA).
However, in the case of AT2019aalc, the large-amplitude variability concluded from the VLASS observations and the EVN detection of a high brightness temperature ( K) feature at GHz clearly imply AGN-like activity over a star formation scenario at lower frequencies. Therefore, our results indicate that the radio spectrum of AT2019aalc is strongly dominated by AGN activity below GHz as well, even if a contribution from star formation, especially at GHz, is probable.
A newly ejected component can explain the high-frequency excess, whilst the low-frequency part originates from optically thin synchrotron emission of the AGN. Thus, the overall spectrum represents a superposition of past and recent AGN activity, indicative of its recently enhanced, intermittent behavior. 2021FrASS...8..147J states that restarted AGN activity can explain at least some of the radio spectra studied by 2020A&A...636A..64B. Intermittent activity is further supported by theoretical models highlighting that sources with high accretion rates are more prone to this kind of behavior (2009ApJ...698..840C).
@@ 5.4 X-ray properties
The X-ray spectra of AGN at energies above keV have a power-law-like shape. These X-ray photons are believed to be triggered by Compton up-scattering of the accretion disk photons off hot electrons surrounding the disk, in a hot ( K) optically thin corona above the disk (e.g., 2015A&A...582A..40K). Below keV some AGN show an excess in their spectrum referred to as the soft X-ray excess. It can be modeled well by a blackbody model with a best-fit temperature in the range keV. The origin of the excess might be explained via Comptonized disk emission. The reflection of hard X-ray photons from the surface of the disk can also account for the excess (2012MNRAS.420.1848D).
The X-ray spectrum of AT2019aalc exhibits a non-thermal component with a power-law index of and an unusually soft thermal component with a blackbody temperature of eV. Most of the BLSy1s typically have a harder X-ray spectrum with a power-law index of (e.g., 2014A&A...568A.108P). AT2019aalc exhibits properties that are more common in sources characterized by enhanced accretion, e.g., NLSy1s or TDEs which both have typically soft X-ray spectra with (1996A&A...305...53B; 2020SSRv..216...85S).
The soft X-ray flux of Seyfert galaxies was found to be correlated well with the observed fluxes in the coronal line [Fe x] . The coronal lines to X-ray flux ratio is log for both broad and narrow-line Seyfert 1 galaxies (2009MNRAS.397..172G). For AT2019aalc we estimate values of log when optical spectroscopic and X-ray observations were performed nearly simultaneously (due to the rapid X-ray variability, we consider only observations taken on the same day). This suggests that this coronal line is excited by the X-ray flares discussed here instead of past X-ray flares that are thought to power the ECLEs.
The reoccurring optical/UV bumps appearing during the decay of the main optical flares are likely driven by the soft X-ray flares, which lead these outbursts only by a couple of weeks. The soft X-ray excess appears to be more dominant during these flares. Interestingly, the soft X-ray flares show hints of quasi-periodicity of days. A few Repeating Nuclear Transients (RNTs) were discovered in the past few years: Quasi-Periodic Eruptions (QPEs) exhibit soft X-ray flares on time-scales of hours to days, while partial TDEs exhibit such flares on timescales of months to years (e.g., qpe2; qpe1; qpe3; qizqpe). We note that another BFF, AT2020afhd, shows hints of quasi-periodic soft X-ray flares according to its Swift/XRT monitoring. This source is included in the The Living Swift XRT Point Source Catalogue (LSXPS: 2023MNRAS.518..174E) as LSXPS J031335.6020905. The potential connection between BFFs and QPEs should be explored in future studies.